OpenStudy (anonymous):

Not sure how to solve this 10,000(1+0.045/12)^12*15

OpenStudy (anonymous):

First to (1+0.045/12)^(12*15)

OpenStudy (anonymous):

1+(0.045/12) = 1+0.00375 = 1.00375. 1.00375^(12*15) = 1.00375^(180) = 1.961555 10000(1.961155) = 19615.5....

OpenStudy (anonymous):

Gosh I love I swear.. I was doing the order of operations wrong I think.